OpenStudy (anonymous):

no.7 and no.8 my last question to my test h.w given that f(x) = 2x^2 - 3x + 4 and g (x) = 3x^2 + 4x - 5, find: 1. (f+g)(-2) 2. (f-g)(-3) 3. (f . g) (2/3) 4. (f/g) (- 4/5) 5. (f o g) (3) 6. (gof)(-4) 7.(f o f) ( 3/2) 8. ( g o g) (-1/2) pls answer 7 and 8 i want to sleep now =)

OpenStudy (anonymous):

7. ( fof)(x)=f(f(x))=f(2x^2 - 3x + 4)=2(2x^2 - 3x + 4)^2-3(2x^2 - 3x + 4)+4 take x=3/2 n simplify

OpenStudy (amistre64):

8. ( g o g) (-1/2) g(g(x)) at x = -1/2 g(-1/2) = 3(-1/2)^2 + 4(-1/2) - 5 = 3(1/4) - 4/2 - 5 = 3/4 - 4/2 - 10/2 = 3/4 - 8/4 - 20/4 = 3/4 - 28/20 = -25/20 g(-25/20) now has to be solved to get the answer

OpenStudy (amistre64):

use a calculator if you understand the concept; otherwise, do it out by hand :)

OpenStudy (anonymous):

thank you sir